OpenStudy (anonymous):

Help! Pic below I give medals :) m arc AC=100deg. m∠ADC=25deg. and C is the point of tangency for DE. Find: m∠ACE m arc BC and m arc AFB

OpenStudy (anonymous):



OpenStudy (mathstudent55):

The measure of an angle formed by a tangent and a secant that intersect outside the circle is half the difference of the arcs. In other words, m<D = (1/2)(m arc AC - m arc BC) Since you know m<D and m arc AC, you can easily get m arc BC. Once you have m arc BC, you can find m arc AFB The measure of an angle formed by a tangent and a chord and whose vertex is a point on the circle is half the measure of the intercepted arc. That means m<ACE = (1/2)m arc AC

OpenStudy (anonymous):

25=m<BC

OpenStudy (anonymous):

m<ACE=50

OpenStudy (mathstudent55):

m<ADC = 25 = m<D m arc AC = 100 m<D = (1/2)(m arc AC - m arc BC) 25 = (1/2)(100 - m arc BC) 50 = 100 - m arc BC -50 = - m arc BC m arc BC = 50

OpenStudy (anonymous):

so how do I find m∠ACE and m arc AFB

OpenStudy (mathstudent55):

You got m<ACE correct. m<ACE = (1/2)m arc AC m<ACE = (1/2)(100) m<ACE = 50

OpenStudy (anonymous):

yes I got that I forgot to simplify

OpenStudy (mathstudent55):

A full circle is an arc of 360 degrees. If m arc AC = 100, and m arc BC = 50, all that's left of the full circle is m arc AFB, so subtract 360 - 100 - 50.